1.      The speed of the bullet (m = 50g) and the speed of an electron ( m = 9.1 * 10-28g) are measured to be the same, namely 300 m/s, with an uncertainty of 0.01%. With what fundamental accuracy could we have located the position of each, if the position is measured simultaneously with the speed in the same expt.?



1
Expert's answer
2022-05-02T13:43:18-0400

"p_ix_i=m_ivx_i=k,\implies x_i=\frac k{m_1v},"

"x_1=6.7\cdot 10^{-6},"

"x_2=3 .7\cdot 10^{23}."